You are interesting in a buying a gift that is shaped like a cube for your fruend.However,you only have a 5 square yards of wrapping paper.the formula is s=√A-6 where s is the edge length,and A is the surface area.What is the longest side length of a cube you could buy based on how much wrapping paper you have?round your answer to the nearest tenth of a yard?


1
Expert's answer
2021-02-24T06:53:19-0500

"\displaystyle
s = \sqrt{\frac{A}{6}} \\
s^2 =\frac{A}{6} \\
A = 6s^2 \\

6s^2 \leq 5 \\

6s^2 - 5 \leq 0\\

s^2 - \frac{5}{6} \leq 0 \\

\left(s - \sqrt{\frac{5}{6}}\right)\left(s - \sqrt{\frac{5}{6}}\right) \leq 0 \\

-\sqrt{\frac{5}{6}} \leq s \leq \sqrt{\frac{5}{6}} \\

\therefore \textsf{The longest side length of a cube you}\\
\textsf{could buy based on how much wrapping}\\
\textsf{paper you have is}\,\,\, 0.9\,\,\textsf{yards}\,\, (\textsf{to nearest tenth})."